# Why Do Open-Source LLMs Struggle with Data Analysis? A Systematic Empirical Study

### v4

@@ 4.3 Strategic Planning
To quantify diversity, we employed a three-stage procedure to classify each question into distinct domains. The detailed methods and descriptions of the categories can be found in Appendix B. From the complete medium turn dataset, we select 2,220 examples under two settings: (1) a natural distribution that reflects the original domain frequencies, and (2) a balanced distribution that down-samples dominant domains while retaining all examples from underrepresented ones. Both settings preserve the overall difficulty distribution of the dataset. Results in Table 7 show minimal performance differences between the two settings, indicating that domain diversity alone does not significantly influence model performance in this context. These findings suggest that the effectiveness of a model is not solely determined by the diversity of problems it encounters during training. Rather, the diversity and richness of reasoning strategies—such as the depth of reasoning processes and the complexity of logical steps—appear to play a more influential role in shaping a model’s performance.
@@ 6 Related Work
To adapt LLMs to complex reasoning tasks, recent work has explored three main approaches: prompt engineering, supervised fine-tuning (SFT), and reinforcement learning (RL) (Wang et al. 2024a; Xi et al. 2025; Guo et al. 2024b; Wang et al. 2024b; Hu et al. 2025; Wang et al. 2025a). Prompt-based methods (Zhao et al. 2024; Yao et al. 2023; Wang et al. 2023) improve reasoning performance by reformulating inputs to better elicit the model’s latent capabilities. SFT-based methods (Chen et al. 2023; Zeng et al. 2024; Yin et al. 2024; Qiao et al. 2024; Song et al. 2024) provide task adaptation through labeled data, and have been applied in settings such as S1 (Muennighoff et al. 2025), which introduces "budget forcing" to control reasoning steps, and LIMO (Ye et al. 2025), which demonstrates that complex mathematical abilities can emerge from carefully curated examples. RL-based approaches further align model behavior with human preferences and problem-solving strategies. Reinforcement learning further enhances reasoning capabilities by optimizing multi-step decision-making. For example, DeepSeek-R1 (DeepSeek-AI et al. 2025) leverages Group Relative Policy Optimization (GRPO) to train LLMs on complex reasoning tasks. RAGEN (Wang et al. 2025b) introduces a modular RL framework that supports stable multi-turn learning via the StarPo architecture. In addition, ReSearch (Chen et al. 2025) integrates RL with retrieval-augmented generation (RAG), treating search operations as part of the reasoning process and learning when and how to retrieve external information.
@@ Acknowledgement
Appendix B Problem Diversity
To quantify diversity, we employed a three-stage procedure to classify each question into distinct domains. In the first stage, we asked the model to classify a small sample of questions without any constraints; Based on its raw outputs, we manually merged semantically similar concepts and split those that were overly broad, yielding a preliminary set of reference categories. In the second stage, we iteratively refined this taxonomy: the model classified each question according to the current reference set but was permitted to propose new categories for outliers, which we then reviewed and either incorporated into or merged with the existing categories. This cycle continued until no further category labels emerged. In the third stage, we constrained the model to assign every question exclusively to one of the finalized categories. Applying this pipeline to our full dataset yielded 10 distinct domains.
@@ Appendix B Problem Diversity
Hyperparameter 7B 14B
batch size 32 32
batch size per device 1 1
gradient accumulation steps 8 4
learning rate $5.0\times 10^{-6}$ $1.0\times 10^{-5}$
epochs 3 3
warmup ratio 0.1 0.1
bf16 true true
Table 10: Detailed training hyperparameters.
Appendix C Data Synthesis Process and Training Details.
The data synthesis pipeline refines responses generated by LLMs to ensure high-quality outputs, as detailed in Algorithm 1. Additionally, we provide some of the prompts we use in our experiments, as shown in Figures  78, which serve as a reference.
@@
10:   if $\text{IsLowComplexity}(r_{\text{clean}})$ or $\text{IsMediumTurn}(r_{\text{clean}})$ then

### v5

@@ Abstract
Codehttps://github.com/zjunlp/DataMind
Extended versionhttps://arxiv.org/abs/2506.19794
@@ 4.3 Strategic Planning
To quantify diversity, we employed a three-stage procedure to classify each question into distinct domains. The detailed methods and descriptions of the categories can be found in Appendix B. From the complete medium turn dataset, we select 2,220 examples under two settings: (1) a natural distribution that reflects the original domain frequencies, and (2) a balanced distribution that down-samples dominant domains while retaining all examples from underrepresented ones. Both settings preserve the overall difficulty distribution of the dataset. Results in Table 7 show minimal performance differences between the two settings, indicating that domain diversity alone does not significantly influence model performance in this context. These findings suggest that the effectiveness of a model is not solely determined by the diversity of problems it encounters during training. Rather, the diversity and richness of reasoning strategies, such as the depth of reasoning processes and the complexity of logical steps, appear to play a more influential role in shaping a model’s performance.
@@ 6 Related Work
To adapt LLMs to complex reasoning tasks, recent work has explored three main approaches: prompt engineering, supervised fine-tuning (SFT), and reinforcement learning (RL) (Wang et al. 2024a; Xi et al. 2025; Guo et al. 2024b; Wang et al. 2024b; Hu et al. 2025; Wang et al. 2025a). Prompt-based methods (Zhao et al. 2024; Yao et al. 2023; Wang et al. 2023) improve reasoning performance by reformulating inputs to better elicit the model’s latent capabilities. SFT-based methods (Chen et al. 2023; Zeng et al. 2024; Yin et al. 2024; Qiao et al. 2024; Zhu et al. 2025) provide task adaptation through labeled data, and have been applied in settings such as S1 (Muennighoff et al. 2025), which introduces "budget forcing" to control reasoning steps, and LIMO (Ye et al. 2025), which demonstrates that complex mathematical abilities can emerge from carefully curated examples. RL-based approaches further align model behavior with human preferences and problem-solving strategies. Reinforcement learning further enhances reasoning capabilities by optimizing multi-step decision-making. For example, DeepSeek-R1 (DeepSeek-AI et al. 2025) leverages Group Relative Policy Optimization (GRPO) to train LLMs on complex reasoning tasks. RAGEN (Wang et al. 2025b) introduces a modular RL framework that supports stable multi-turn learning via the StarPo architecture. In addition, ReSearch (Chen et al. 2025) integrates RL with retrieval-augmented generation (RAG), treating search operations as part of the reasoning process and learning when and how to retrieve external information.
@@ Acknowledgement
Appendix B Problem Diversity.
To quantify diversity, we employed a three-stage procedure to classify each question into distinct domains. In the first stage, we asked the model to classify a small sample of questions without any constraints; Based on its raw outputs, we manually merged semantically similar concepts and split those that were overly broad, yielding a preliminary set of reference categories. In the second stage, we iteratively refined this taxonomy: the model classified each question according to the current reference set but was permitted to propose new categories for outliers, which we then reviewed and either incorporated into or merged with the existing categories. This cycle continued until no further category labels emerged. In the third stage, we constrained the model to assign every question exclusively to one of the finalized categories. Applying this pipeline to our full dataset yielded 10 distinct domains as shown in Figure 6.
@@
Appendix C Data Synthesis Details
@@
10:   if $\text{IsLowComplexity}(r_{\text{clean}})$ or $\text{NotMediumTurn}(r_{\text{clean}})$ then
@@
Our data synthesis pipeline (see Algorithm 1) transforms raw queries into high-quality instances for Supervised Fine-Tuning. The process consists of three core stages:
  1. 1.
    Prompt-Based Answer Generation: For each input question $q\in\mathcal{Q}$, we use a base generation prompt (Figure 7) and a non-zero temperature (e.g., $T=0.6$) to generate a set of diverse candidate responses $R_{q}=\{r_{1},\dots,r_{k}\}$, where we set $k=3$ in our experiments.
  2. 2.
    Targeted Instance Selection: Next, we conduct a multi-faceted filtering process where each candidate response must pass the following sequential criteria to be selected:
    • Correctness Verification: First, as a crucial preliminary step, we execute the code within each candidate response to verify its correctness against a ground-truth solution. Only demonstrably correct responses proceed to the subsequent filters.
    • Format Conformance: Next, we apply a set of rules to each verified response to standardize its structure, producing a cleaned version, $r_{\text{clean}}$. Responses that fail to conform to our required format (e.g., missing or malformed code blocks) are discarded.
    • Length and Difficulty Filtering: Finally, the cleaned response $r_{\text{clean}}$ is subjected to two concurrent checks. We retain only those spanning 4–5 turns (medium length) and discard any solvable by a less capable baseline model (Qwen2.5-7B-Instruct) to ensure medium-to-high difficulty.
    Only the responses that pass all these filters are passed to the next stage.
  3. 3.
    Reasoning-Driven Data Enrichment: For each filtered response $r_{\text{clean}}$, we first extract its detailed reasoning (reason_chain) and then condense it into a concise summary, reason_summary. This is accomplished using a dedicated summarization prompt (Figure 8). The final instance, $r_{\text{final}}$, is formed by inserting this summary into the response, explicitly capturing the underlying problem-solving strategy.
This pipeline yields a final dataset of 2.8k instances used for our fine-tuning experiments.
Hyperparameter 7B 14B
batch size 32 32
batch size per device 1 1
gradient accumulation steps 8 4
learning rate $5.0\times 10^{-6}$ $1.0\times 10^{-5}$
epochs 3 3
warmup ratio 0.1 0.1
bf16 true true
Table 10: Detailed training hyperparameters.
Training Details.